### v2

@@ 3.4.2 Tuning $\rho$
We now outline a data-adaptive procedure to select $\rho$, where the goal is to prevent information from the large confounded observational study from swamping that of the smaller unconfounded RCT. Our proposal is based on cross-validation to minimize the Root Mean Squared Error (RMSE) on weighted held-out RCT patients only, where the weighting tailors $\rho$ for extrapolation. Specifically, use 5-fold cross-validation, as the RCT sample size is usually small, and we consider $\rho\in[0.0,0.1,\cdots,1.0]$ on a grid. In particular, we minimize the following objective:
@@ 3.4.3 Extension to CATE Estimation
When both treatment and control arms are available, we fit two separate rank-2 ICMs to model $(f_{0}^{\text{e}},f_{0}^{\text{o}})$ and $(f_{1}^{\text{e}},f_{1}^{\text{o}})$, focusing on sharing information between the experimental and observational datasets. This yields independent posteriors for $f_{0}^{\text{e}}$ and $f_{1}^{\text{e}}$, which we combine to obtain the posterior of $\tau(\mathbf{x})$ by differencing the means and summing the variances. We allow distinct kernel hyperparameters for the two ICMs while using a common $\rho$, chosen to minimise the average $\mathcal{L}(\rho)$ across treatment groups. This setup corresponds to a T-learner approach (Künzel et al., 2019).
@@ 4 Experiments
We investigated Causal-ICM’s performance through multiple simulation studies and a comprehensive analysis of Real-World Data (RWD) sourced from the Tennessee STAR study (Achilles et al., 2008). The code will be made available online. The optimal value of $\rho$ is chosen data-adaptively in all cases (Section 3.4.2). Model evaluation and comparison were based on the RMSE, averaged over the covariate distribution of the observational study.
We compared Causal-ICM against several benchmarks: (i) a T-learner with GPs regressors trained separately on each study, (ii) the two-step method of Kallus et al. (2018), which debiases observational data via a low-complexity bias function with GPs as base learners, (iii) the integrative estimator of Yang et al. (2022b), which assumes a non-linear CATE and linear confounding effect, (iv) the power likelihood method of Lin and Evans (2023), and (v) the test-then-pool approach of Yang et al. (2022a). For Yang et al. (2022b), we used the default settings recommended in their paper, while for Lin and Evans (2023), we selected the power parameter $\eta$ by maximizing the Expected Log Pointwise Predictive Density (ELPD), which in both cases resulted in $\eta=0$ (no pooling). Finally, we assessed the coverage of Causal-ICM’s CATE credible intervals across simulation scenarios, comparing them with those from the T-learner approaches.
We used GPy (2012) to train the ICMs, with $\rho$ tuned as described in Section 3.4.2 and remaining hyperparameters estimated via marginal likelihood maximisation. All experiments were run on a single Macbook Pro (M2) with 16GB RAM.
@@ 4.1 Simulation Studies
We showcase the performance of Causal-ICM in the main text using two distinct simulation studies. Further simulation studies, both univariate and multivariate, are detailed in Appendix A.8, illustrating the competitive performance of Causal-ICM in a wide range of settings, including both linear and non-linear scenarios for the potential outcomes, CATE and confounding function. In all simulations, the sample size is roughly between 200-300 for the RCT and 1000 for the observational study.
@@
In the first simulation setting, we assume a continuous baseline covariate $X\sim\text{Unif}\left[-2,2\right]$. The values of $X$ control the trial participation probability, where $S\sim\text{Bernoulli}\left(p_{S}\right)$ for $p_{S}={\text{exp}\left(-3-3X\right)}/{(1+\text{exp}\left(-3-3X\right)})$. Here, patients with smaller $X$ values have higher probability of being assigned in the trial. After selecting all trial participants, we assign them randomly to treatment with $A\sim\text{Bernoulli}\left(0.5\right)$. The observed outcomes are generated as $Y=A\tau\left(X\right)+X+\epsilon$ for $A\in\{0,1\}$, where $\tau\left(X\right)=1+X$ is the CATE and $\epsilon\sim\mathcal{N}\left(0,1\right)$. For the observational study we have $X\sim\text{Unif}\left[-2,2\right]$ and the treatment is generated according to $A\sim\text{Bernoulli}\left(e\left(X\right)\right)$ where $\text{logit}\left(e\left(X\right)\right)=-X$. Similarly to the trial, the observed outcomes are $Y=A\tau\left(X\right)+X+U+\epsilon$ for $A\in\{0,1\}$. Here, $U$ is the hidden confounder which is generated from $U\sim\mathcal{N}\left(\left(2A-1\right)X,1\right)$, yielding a linear confounding effect $\eta(X)=2X$. In the second simulation, we change the potential outcomes, the CATE and the confounding effect to be non-linear as follows: $Y=A\tau(X)+X^{2}-1+U+\epsilon$, where $\tau(X)=1+X+X^{2}$ and $U\sim\mathcal{N}((2A-1)\sin(X-1),1)$, yielding $\eta(X)=2\text{sin}(X-1)$. All other variables remain the same.
@@
Under a linear CATE and a linear unobserved confounding effect, Causal-ICM (with $\rho=0.1$) exhibited comparable performance to the integrative HTE estimator (Yang et al. (2022b)) and the Elastic HTE estimator of Yang et al. (2022a), and superior performance to all other techniques (Fig 1 left). In the non-linear setting (both CATE and unobserved confounding effect), Causal-ICM outperformed all other methods yielding the lowest RMSE (Fig 1 right). These results reflect the limitations of the experimental grounding, the Integrative HTE and the Elastic HTE methods. Each of these are designed for linear treatment or confounding effect, resulting in inflated RMSE values in scenarios where these assumptions do not hold. In the multivariate setting Causal-ICM performed equally good or better than most methods, being outperformed by parametric approaches only when the model was correctly specified (see Appendix G.2 Fig 4).
@@
Regarding uncertainty quantification, our approach generated 95% credible intervals with reasonable, though slightly conservative, coverage close to the nominal level in both simulation settings (Fig 2). The observational-only T-learner had the lowest coverage across the baseline covariate support, which is expected since its estimates are biased but exhibit lower uncertainty due to the larger dataset. This is exactly the behaviour that Causal-ICM is robust to, as formalized in Proposition 3.1: the large sample size of the observational study does not overwhelm the Causal-ICM’s estimates. See Figure 5 in Appendix H for an illustrative example of extrapolation and uncertainty quantification with Causal-ICM in the second simulation setting. Additionally, the Causal-ICM is robust against the kernel choice (see Appendix I Table 3), the observational sample size (see Appendix I Table 4), and the degree of overlap between the two studies (see Appendix I Table 5), yielding similar RMSE values under all these different scenarios. Finally, the choice of $\rho$ plays an important role in controlling the influence of the confounded observational data (see Appendix I Table 2).
@@ 4.2 Real-World Data Analysis
The Tennessee Student/Teacher Achievement Ratio (STAR) (Achilles et al., 2008) randomised experiment initiated in 1985 studied the impact of class size on student outcomes through standardized test scores from first to third grade. We focus on two ‘treatments’, here corresponding to two experimental conditions: small class size and regular class size. We followed a similar approach to Kallus et al. (2018) and used the real data to produce a ‘confounded’ dataset, corresponding to the observational sample, and a smaller unconfounded dataset, which comprised our RCT data. After removing subjects with missing treatment or outcome values, we were left with a randomised sample of 4218 students. Baseline covariates included gender, race, birth month, birth year, free lunch eligibility and teacher ID. To obtain the ‘experimental’ dataset we sampled randomly only from rural or inner city students. To form the confounded observational dataset, we aggregated all control cases not included in the experimental data. Additionally, for those undergoing treatment, we drew a sample with a down-weighting mechanism applied to individuals whose outcomes fell below the $30^{th}$ percentile. After preprocessing, the unconfounded dataset included 422 students, while the confounded dataset included 2593. Further, 379 students were kept as a validation set. We compared Causal-ICM (optimal $\rho=0.4$) with a GP based T-learner trained either on the experimental or the observational data,the experimental grounding method of Kallus et al. (2018) either with Random Forest or GP as base learners and the Integrative HTE method of Yang et al. (2022b). We were unable to include the Elastic HTE estimator (Yang et al. (2023)) or the power likelihood approach (Lin et al. (2025a)), as both methods encountered convergence issues due to the datasets’ high dimensionality and the limited variability in their categorical covariates.
@@
In order to compare Causal-ICM with existing methods, we require a notion of ground truth. Given the randomised nature of the study, we assume that an unbiased estimate of the CATE can be obtained using a doubly robust estimator, which we then treat as the ground truth. To obtain this, we estimated the propensity score and conditional expectation models using the full dataset, prior to splitting it into the confounded and unconfounded samples (Saito and Yasui, 2020). We note that, in the absence of the true CATE, other performance metrics adapted to the counterfactual nature of the evaluation are also available (Alaa and Van Der Schaar, 2019; Boyer et al., 2023).
The results are summarised in Table 1 below. Causal-ICM again performed highly competitively in terms of RMSE, on a par with the experimental grounding method with Random Forests as base learners.
@@
Table 1: RMSE values for the RWD example. Causal-ICM: our method; GP (exp): T-learner with GPs as base learners trained in the unconfounded; GP (obs): T-learner with GPs as base learners trained in the confounded data; Experimental grounding (GP): the method proposed by Kallus et al. (2018) with GPs as base learners; Experimental grounding (RF): the method proposed by Kallus et al. (2018) with Random Forest as base learners; Integrative HTE: the method proposed by Yang et al. (2022b)
@@ 5 Discussion
We have developed Causal-ICM, a rank-2 ICM to combine observational and experimental data to estimate treatment effects for a target population, where the hyperparameter $\rho$ controls the degree of borrowing in an interpretable way. In contrast to existing methods, Causal-ICM is highly flexible and does not make any assumptions about the functional form of the CATE or the effect of unobserved confounding. By providing accurate estimates of heterogeneous treatment effects, Causal-ICM has the potential to inform more personalised medical decision making. Moreover, unlike other data fusion methods, Causal-ICM is able to provide uncertainty quantification. By identifying regions of the population where treatment effects remain highly uncertain, our approach can help guide the design of future trials to improve representativeness and reduce uncertainty to satisfactory levels (Yang et al. (2025)).
Despite the inherent difficulty of extrapolation, Causal-ICM appears to extrapolate well, achieving reliable levels of coverage. In general, it is challenging to ensure frequentist coverage for Bayesian nonparametric methods, which is a limitation of our approach. Another limitation is the sensitivity of uncertainty estimates to the kernel choice and the hyperparameter optimization, which is an intrinsic challenge of GPs. Future research directions for Causal-ICM include extensions to multiple data sources or treatment arms, placing a prior on the parameter $\rho$, and incorporating the propensity score into the model. Investigating partial identification of causal effects, especially in regions of extrapolation, is also important Pfister and Bühlmann (2024). Methodological extensions such as allowing distinct kernels for the latent functions, accommodating non-Gaussian likelihoods, and improving computational scalability via sparse GPs or inducing points represent additional promising directions (see Appendix J).
@@ Appendix A Identification of CATE in the RCT
(tower property)
Appendix B Illustrative example
@@
Appendix C Derivation of Posterior Distribution
@@ Appendix D Interpreting $\rho$
Appendix E Proof of Proposition 3.1
@@
Appendix F Confounding function

### v3

@@ 3.4.2 Tuning $\rho$
We now outline a data-adaptive procedure to select $\rho$, where the goal is to prevent information from the large confounded observational study from swamping that of the smaller unconfounded RCT. Our proposal is based on cross-validation to minimize the Root Mean Squared Error (RMSE) on weighted held-out RCT patients only, where the weighting tailors $\rho$ for extrapolation. Specifically, we use 5-fold cross-validation, as the RCT sample size is usually small, and we consider $\rho\in[0.0,0.1,\cdots,1.0]$ on a grid. Our objective is:
@@ 3.4.3 Extension to CATE Estimation
When both treatment and control arms are available, we fit two separate rank-2 ICMs to model $(f_{0}^{\text{e}},f_{0}^{\text{o}})$ and $(f_{1}^{\text{e}},f_{1}^{\text{o}})$, focusing on sharing information between the experimental and observational datasets. This yields independent posteriors for $f_{0}^{\text{e}}$ and $f_{1}^{\text{e}}$, which we combine to obtain the posterior of $\tau(\mathbf{x})$ by differencing the means and summing the variances. We allow distinct kernel hyperparameters for the two ICMs while using a common $\rho$, chosen to minimise the average $\mathcal{L}(\rho)$ across treatment groups. This setup corresponds to a T-learner approach (Künzel et al., 2019), and avoids imposing strong assumptions on the bias structure across treatment arms and preserves theoretical tractability. An alternative approach would be to also couple across treatment arms; however, in preliminary experiments not presented here, we found that such strongly coupled models resulted in over-shrinkage of posterior variance and reduced coverage.
@@ 4 Experiments
We investigated Causal-ICM’s performance through multiple simulation studies and a comprehensive analysis of Real-World Data (RWD) sourced from the Tennessee STAR study (Achilles et al., 2008). The code is available online [^1] https://github.com/EvanDimitriou/CausalICM. The optimal value of $\rho$ is chosen data-adaptively in all cases (Section 3.4.2). Model evaluation and comparison were based on the RMSE, averaged over the covariate distribution of the observational study.
We compared Causal-ICM against several benchmarks: (i) a T-learner with GPs regressors trained separately on each study, (ii) the two-step method of Kallus et al. (2018), which debiases observational data via a low-complexity bias function with GPs as base learners, (iii) the integrative estimator of Yang et al. (2025b), which assumes a non-linear CATE and linear confounding effect, (iv) the power likelihood method of Lin et al. (2025), and (v) the test-then-pool approach of Yang et al. (2023). For Yang et al. (2025b), we used the default settings recommended in their paper, while for Lin et al. (2025), we selected the power parameter $\eta$ by maximizing the Expected Log Pointwise Predictive Density (ELPD), which in both cases resulted in $\eta=0$ (no pooling). Finally, we assessed the coverage of Causal-ICM’s CATE credible intervals across simulation scenarios, comparing them with those from the T-learner approaches.
We used GPy (2012) to train the ICMs, with $\rho$ tuned as described in Section 3.4.2 and remaining hyperparameters estimated via marginal likelihood maximisation. All experiments were conducted on a MacBook Pro (2022) equipped with an Apple M2 chip, 8 CPU cores, 10 GPU cores, and 16 GB of memory, running macOS 26.2. Runtime comparisons indicate that Causal-ICM incurs a moderate computational overhead relative to single-study GP baselines, but remains competitive with other data-fusion approaches (see Appendix K).
@@ 4.1 Simulation Studies
We showcase the performance of Causal-ICM in the main text using two distinct univariate simulation studies. Multivariate simulation studies, are detailed in Appendix H, illustrating the competitive performance of Causal-ICM in higher dimensions. In all simulations, the sample size is roughly between 200-300 for the RCT and 1000 for the observational study.
@@
In the first simulation setting, we assume a continuous baseline covariate $X\sim\text{Unif}\left[-2,2\right]$. The values of $X$ control the trial participation probability, where $S\sim\text{Bernoulli}\left(p_{S}\right)$ for $p_{S}={\text{exp}\left(-3-3X\right)}/{(1+\text{exp}\left(-3-3X\right)})$. Here, patients with smaller $X$ values have higher probability of being assigned in the trial. Trial participants are randomly assigned treatment with $A\sim\text{Bernoulli}\left(0.5\right)$. The observed outcomes are generated as $Y=A\tau\left(X\right)+X+\epsilon$ for $A\in\{0,1\}$, where $\tau\left(X\right)=1+X$ is the CATE and $\epsilon\sim\mathcal{N}\left(0,1\right)$. For the observational study we have $X\sim\text{Unif}\left[-2,2\right]$, with the treatment generated according to $A\sim\text{Bernoulli}\left(e\left(X\right)\right)$ where $\text{logit}\left(e\left(X\right)\right)=-X$. The observed outcomes are $Y=A\tau\left(X\right)+X+U+\epsilon$ for $A\in\{0,1\}$. Here, the hidden confounder is generated from $U\sim\mathcal{N}\left(\left(2A-1\right)X,1\right)$, yielding a linear confounding effect $\eta(X)=2X$. In the second simulation, we have non-linear potential outcomes, CATE and confounding effect, as follows: $Y=A\tau(X)+X^{2}-1+U+\epsilon$, where $\tau(X)=1+X+X^{2}$ and $U\sim\mathcal{N}((2A-1)\sin(X-1),1)$, yielding $\eta(X)=2\text{sin}(X-1)$. All other variables remain the same.
@@
Under a linear CATE and a linear unobserved confounding effect, Causal-ICM with $\rho=0.1$ exhibited comparable performance to the integrative HTE (Yang et al. (2025b)) and Elastic HTE estimators (Yang et al. (2023)), and outperformed all other techniques (Fig 1 left). In the non-linear setting (both CATE and unobserved confounding effect), Causal-ICM outperformed all other methods yielding the lowest RMSE (Fig 1 right). These results reflect the limitations of the experimental grounding, the Integrative HTE and the Elastic HTE methods. Each of these are designed for linear treatment or confounding effect, resulting in inflated RMSE values in scenarios where these assumptions do not hold.
@@
Regarding uncertainty quantification, our approach generated 95% credible intervals with reasonable, though slightly conservative, coverage close to the nominal level in both simulation settings (Fig. 2). This conservativeness is consistent with the structure of the model: posterior uncertainty reflects not only sampling variability, but also uncertainty induced by partial cross-domain borrowing and the possibility of residual observational bias. The observational-only T-learner had the lowest coverage across the baseline covariate support, which is expected in the presence of unmeasured confounding: as sample size grows, its estimates can remain biased while posterior uncertainty shrinks, leading to overconfident credible intervals that fail to cover the true CATE. This is exactly the behaviour that Causal-ICM is robust to, as formalized in Proposition 3.1: the large sample size of the observational study does not overwhelm the Causal-ICM estimates. See Figure 5 in Appendix I for an illustrative example of extrapolation and uncertainty quantification. Additional sensitivity analysis regarding imbalance of study sample sizes, performance within and outside the RCT support, kernel choice and the degree of overlap between the two studies and the choice of $\rho$ is provided in Appendix J.
@@ 4.2 Real-World Data Analysis
The Tennessee Student/Teacher Achievement Ratio (STAR) (Achilles et al., 2008) randomised experiment initiated in 1985 studied the impact of class size on student outcomes through standardized test scores from first to third grade. We focus on two ‘treatments’, here corresponding to two experimental conditions: small class size and regular class size. We followed a similar approach to Kallus et al. (2018) and used the real data to produce a ‘confounded’ dataset, corresponding to the observational sample, and a smaller unconfounded dataset, which comprised our RCT data. After removing subjects with missing treatment or outcome values, we were left with a randomised sample of 4218 students. Baseline covariates included gender, race, birth month, birth year, free lunch eligibility and teacher ID. To obtain the ‘experimental’ dataset we sampled randomly only from rural or inner city students. To form the confounded observational dataset, we aggregated all control cases not included in the experimental data. Additionally, for those undergoing treatment, we drew a sample with a down-weighting mechanism applied to individuals whose outcomes fell below the $30^{th}$ percentile. After preprocessing, the unconfounded dataset included 422 students, while the confounded dataset included 2593. Further, 379 students were kept as a validation set. We compared Causal-ICM (optimal $\rho=0.4$) with a GP based T-learner trained either on the experimental or the observational data,the experimental grounding method of Kallus et al. (2018) either with Random Forest or GP as base learners and the Integrative HTE method of Yang et al. (2025b). We were unable to include the Elastic HTE estimator (Yang et al. (2023)) or the power likelihood approach (Lin et al. (2025)), as both methods encountered convergence issues due to the datasets’ high dimensionality and the limited variability in their categorical covariates.
@@
In order to compare Causal-ICM with existing methods, we require a notion of ground truth. Given the randomised nature of the study, we assume that an unbiased estimate of the CATE can be obtained using a doubly robust estimator, which we then treat as the ground truth. To obtain this, we estimated the propensity score and conditional expectation models using the full dataset, prior to splitting it into the confounded and unconfounded samples (Saito and Yasui, 2020). We note that, in the absence of the true CATE, other performance metrics adapted to the counterfactual nature of the evaluation are also available (Alaa and Van Der Schaar, 2019; Boyer et al., 2023). The results are summarised in Table 1. Causal-ICM again performed highly competitively in terms of RMSE, on a par with the experimental grounding method with Random Forests as base learners.
@@
Table 1: RMSE values for the RWD example. Causal-ICM: our method; GP (exp): T-learner with GPs as base learners trained in the unconfounded; GP (obs): T-learner with GPs as base learners trained in the confounded data; Experimental grounding (GP): the method proposed by Kallus et al. (2018) with GPs as base learners; Experimental grounding (RF): the method proposed by Kallus et al. (2018) with Random Forest as base learners; Integrative HTE: the method proposed by Yang et al. (2025b).
@@ 5 Discussion
We have developed Causal-ICM, a rank-2 ICM to combine observational and experimental data to estimate treatment effects for a target population, where the hyperparameter $\rho$ controls the degree of borrowing in an interpretable way. In contrast to existing methods, Causal-ICM is highly flexible and does not make assumptions about the functional form of the CATE or the effect of unobserved confounding. By construction, the method trades a small amount of potential bias from the observational source against gains in precision and external validity, while anchoring inference to the RCT. This trade-off is further supported by our additional experiments, which show that Causal-ICM remains stable under increasing observational sample sizes and continues to perform competitively even when alternative data-fusion approaches deteriorate. By providing accurate estimates of HTEs together with uncertainty quantification, Causal-ICM has the potential to inform more personalised medical decision making. By identifying regions of the population where treatment effects remain highly uncertain, our approach can help guide the design of future trials to improve representativeness and reduce uncertainty to satisfactory levels (Yang et al. (2025a)).
Despite the inherent difficulty of extrapolation, Causal-ICM performs well in these settings, as evidenced by our additional analysis separating performance within and outside the support of the randomized trial, where the gains are particularly pronounced in regions lacking experimental data. Causal-ICM also achieves reasonable, if slightly conservative, credible intervals. In general, it is challenging to ensure frequentist coverage for Bayesian nonparametric methods. The slightly conservative empirical coverage of Causal-ICM can be explained by the model hedging against disagreement between the experimental and observational signals, with additional uncertainty induced by partial cross-domain borrowing. A related limitation of our approach is the sensitivity of uncertainty estimates to the kernel choice and the hyperparameter optimization, itself an intrinsic challenge of GPs.
The flexibility of the (multi-task) GP framework underlying Causal-ICM opens several avenues for future work. In particular, the current model uses shared kernel hyperparameters across latent functions but more flexible alternatives, such as the Linear Model of Coregionalization (LCM; Alvarez et al. (2012)), could allow task-specific structure. Other possible extensions include incorporating more than two data sources or treatment arms, and placing a prior on the borrowing parameter $\rho$. More broadly, extending the framework to allow non-Gaussian likelihoods (e.g. for binary or heavy-tailed outcomes) is of significant practical interest, but will likely require non-trivial posterior approximations.
@@ Appendix A Identification of CATE in the RCT
(tower property and $P(X_{S}|X_{\tau})=P(X_{S}|X_{\tau},S=e)$)
Appendix B Discussion of Assumptions
Assumption A4 is weaker than requiring the RCT to be a random sample from the target population: conditional on $\mathbf{X}_{\tau}$, it only requires the relevant distribution of selection variables to agree between the trial and the target population. Note that this is strictly weaker than assuming that the RCT is a random sample from the target population and instead allows for some level of distributional shift. Formally, this means that the CATE in the target population can be identified as
$\tau(X_{\tau})=\mathbb{E}[Y^{1}-Y^{0}|X_{\tau}]=\mathbb{E}_{X_{S}|X_{\tau}}[\mathbb{E}[Y^{1}-Y^{0}|X_{\tau},X_{S},S=e]]$
Thus averaging over $X_{S}$ given $X_{\tau}$ in the RCT correctly recovers the CATE. Importantly, this does not require the marginal distribution of $X_{S}$ to be identical across studies, and distribution shift on $X_{\tau}$ (marginally) is allowed.
Although the CATE can be identified within the RCT population under assumptions A1-A5, this highlights a fundamental limitation of relying on a single data source. The identified estimand, $\tau(X_{\tau})=\omega^{(e)}(X_{\tau})$, pertains only to the RCT population and therefore reflects the causal effect among individuals who were eligible for and participated in the trial. In many applied settings, however, the primary scientific or policy objective is to infer treatment effects in a broader target population, which may be better represented by an observational study. Assumption A6 formalises the conditions under which such transport is possible: within levels of $X_{\tau}$, the distribution of $X_{S}$ is invariant between the RCT and the target population, while still allowing for marginal differences in $X_{S}$ across the two populations.
Appendix C Illustrative example
@@
Appendix D Derivation of Posterior Distribution
@@ Appendix E Interpreting $\rho$
Appendix F Proof of Proposition 3.1
@@
Appendix G Confounding function